# Towards Sustainable Universal Deepfake Detection with Frequency-Domain Masking

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOMMccs: Computing methodologies Biometrics
@@ 1. Introduction
The proliferation of increasingly convincing synthetic images, facilitated by generative AI, poses significant challenges across multiple sectors, including cybersecurity, digital forensics, and public discourse (ojha2023fakedetect; Gragnaniello2021; DBLP:conf/eccv/ChaiBLI20; Abdollahzadeh2023ASO). These AI-generated images could be misused as deepfakes for malicious purposes, such as disinformation. Because these fake images can look very real, they can trick people and systems easily. Detection of deepfakes is an important problem that has attracted significant attention.
@@
Universal deepfake detection. Early research in deepfake detection primarily focused on identifying synthetic images generated by specific types of generative models, such as Generative Adversarial Networks (GANs) (mirsky2020deepfake). However, with the rapid advancement of generative methods, including the rise of diffusion models (rombach2021diffusion), the need for universal deepfake detection has become more prominent. This approach aims to detect deepfakes across a wide variety of generative AI models, including those that were not part of the training data for the detection system. Such a shift highlights the necessity for deepfake detection systems to exhibit robust generalization capabilities. In the early stages, Wang et al. (wang2019cnngenerated) explored the use of post-processing and data augmentation techniques to detect synthetic images, focusing particularly on those generated by GANs. Later, Chandrasegaran et al. (Chandrasegaran_2022_ECCV) introduced the concept of transferable forensic features, which are generalizable across different generative models, thus improving the detector’s ability to handle unseen models. More recent approaches, such as those by Ojha et al. (ojha2023fakedetect), leverage the feature space of large pretrained models, showing strong performance across unseen generators, although they depend heavily on access to large-scale pretrained models. A notable advancement by Tan et al. (Tan2024RethinkingTU) investigated the impact of architectural artifacts, specifically those introduced by upsampling operations in generative models, and proposed detecting structural inconsistencies to improve generalization. Lastly, Chen et al. (Chen2022OSTIG) examined one-shot test-time training to enhance model generalization; however, this technique introduces a trade-off between accuracy and computational efficiency due to the need for additional resources per test sample.
@@
The case for Green AI. Universal deepfake detection, which aims for strong generalization across a variety of scenarios, often depends on large pre-trained models (ojha2023fakedetect) and specialized feature extraction techniques (Tan2024FrequencyAwareDD), both of which demand substantial computational resources. This dependency raises concerns about the scalability of these methods and their negative environmental impact due to the high energy consumption associated with training and inference. In response, the machine learning community is increasingly focusing on the concept of Green AI, which promotes the development of methods capable of achieving competitive performance while minimizing energy usage and computational cost (chen2024gift; Kuo2022GreenLI; Zhu2022RGGIDAR; Zhu2021APixelHopAG; Wei2024AGL). This shift in focus has led to efforts aimed at creating lightweight yet generalizable detection strategies that remain effective even in environments with limited resources. In this regard, frequency-domain representations stand out as an especially promising solution (Coccomini2024DeepfakeDW). These representations preserve important structural and statistical information, typically with fewer dimensions than their spatial counterparts, and require significantly lower computational overhead. Consequently, employing a frequency-based approach offers the potential to enhance both the generalization ability and efficiency of deepfake detection models. Beyond general-purpose detection, specialized domains (e.g., aquaculture) increasingly rely on synthetic data for training and augmentation. However, poor-quality synthetic samples may propagate errors in critical applications—from misdiagnosing fish diseases to inaccurate medical assessments. A robust universal detector must therefore generalize not only across generative architectures but also across diverse application domains with potentially limited training data.
@@
Masked image modeling. In recent years, masked image modeling (MIM) has gained significant attention as an effective self-supervised pre-training technique to enhance the generalization ability of models (DBLP:conf/cvpr/HeCXLDG22; DBLP:journals/corr/abs-2302-02615; DBLP:conf/iclr/0002LZ0OL23; DBLP:conf/nips/HuangCGXZLLX22). This method revolves around pre-training models to predict missing or masked parts of unlabeled data, with reconstruction loss serving as the objective for learning. After the pre-training phase, the trained encoder can be effectively fine-tuned for a variety of downstream tasks, demonstrating strong transferability. He et al. (DBLP:conf/cvpr/HeCXLDG22) demonstrated that training with a masked autoencoder (MAE) allows high-capacity models to achieve state-of-the-art performance in terms of generalization across diverse applications. More recent works, such as those by Li et al. (DBLP:journals/corr/abs-2302-02615), argue that reconstruction-based objectives excel at learning in-distribution representations, making them particularly effective for out-of-distribution detection. Xie et al. (DBLP:conf/iclr/0002LZ0OL23) expanded this approach by incorporating frequency-domain information into the masking process, proposing Masked Frequency Modeling (MFM), where the model learns to predict masked frequency components rather than spatial patches. This enables models to capture structured global image priors more efficiently. In the context of deepfake detection, Das et al. (Das2023UnmaskingDM) leveraged spatiotemporal transformers and masked autoencoding, showing that incorporating both spatial and temporal inconsistencies improves generalization across datasets.
@@
We analyzed both random masking and geometric transformations for universal deepfake detection. Our results suggest that frequency masking is more effective in generalizing detection across diverse generative AI approaches. Our finding is consistent with a recent study by Corvi et al. (Corvi_2023_CVPR), which identifies frequency artifacts in GAN- and diffusion-based synthetic images. Different from (Corvi_2023_CVPR), our main contribution is a new training method that improves detection accuracy and robustness via frequency-domain masking. We further demonstrate that this method is resilient to model compression, i.e., structured pruning, making it attractive for real-world, resource-constrained deployments.
We remark that most existing detectors focus on spatial-domain artifacts (ojha2023fakedetect; Gragnaniello2021; wang2019cnngenerated). Our contributions are summarized as follows:
@@ 2.1. Universal Deepfake Detection
Early research in universal deepfake detection focused on identifying common artifacts in synthetic images. Zhang et al. (Zhang2019DetectingAS) proposed simulating GAN artifacts to train frequency-domain classifiers capable of generalizing across unseen models. Wang et al. (wang2019cnngenerated) discovered that CNN-generated images exhibit systematic artifacts detectable by CNN-based classifiers, even when the detector is trained on a single generator. This finding suggested shared weaknesses among early GAN models. In contrast, Gragnaniello et al. (Gragnaniello2021) emphasized the growing realism of synthetic content and highlighted the necessity of robust and generalizable detection tools in the face of increasingly sophisticated generative models. Chandrasegaran et al. (Chandrasegaran_2022_ECCV) introduced the idea of transferable forensic features, demonstrating that certain features—particularly color-based ones—can be learned to generalize across models. Ojha et al. (ojha2023fakedetect) proposed leveraging the feature space of large pretrained models for real-vs-fake classification. This method, without explicit model-specific training, was shown to outperform traditional classifiers on unseen generative models. More recently, Tan et al. (Tan2024RethinkingTU) explored architectural artifacts, particularly those introduced by upsampling operations in generative models. They proposed detecting structural inconsistencies such as neighboring pixel relationships (NPR), which yielded high generalization performance on unseen generators. Compared to these approaches, our method introduces masked image modeling as a training-time augmentation that reduces overfitting to generator-specific traces.
@@ 2.2. Frequency-Domain Analysis
Pioneering studies have explored frequency-domain cues as critical features for improving the robustness and generalizability of deepfake detection systems. Frank et al. (Frank2020LeveragingFA) first demonstrated that GAN-generated images often exhibit unnatural frequency artifacts due to upsampling operations, which can be effectively exploited to identify fakes. Qian et al. (Qian2020ThinkingIF) further investigated these frequency-aware clues, showing that subtle forgery traces and compression-induced anomalies manifest more clearly in the frequency domain, especially in low-quality images. Luo et al. (Luo2021GeneralizingFF) addressed the generalization challenge across forgery methods by proposing three modules to extract high-frequency noise patterns, which generalize better than spatial-domain textures. Complementary to this, Le and Woo (Le2021ADDFA) introduced a knowledge distillation framework where a student network learns from high-quality teacher signals in the frequency domain, enabling robust detection of compressed and degraded images. Jeong et al. (Jeong2022FrePGANRD) recognized the risk of overfitting to frequency-specific artifacts and proposed FrePGAN, which transitions from frequency-focused learning to incorporating image-level irregularities, thus balancing generalization across known and unknown generators. Tian et al. (Tian2023FrequencyAwareAF) combined RGB and frequency representations using attention-based fusion, showing that this dual-modality approach alleviates overfitting and enhances generalization. Most recently, Tan et al. (Tan2024FrequencyAwareDD) developed a frequency-aware architecture that targets high-frequency patterns across spatial and channel dimensions to improve generalizability while maintaining low model complexity. Li et al. (Li2024FreqBlenderED) introduced a data generation strategy that creates pseudo-fake images by blending frequency-domain features, enriching training data to better capture generalizable forgery traits. Distinctly, our approach employs frequency-domain masking as a supervised augmentation method that promotes learning of features resilient to domain shift.
@@ 2.3. Masked Image Modeling
Masked image modeling (MIM) has emerged as a powerful self-supervised learning paradigm in computer vision, with increasing relevance to generalization and robustness. He et al. (DBLP:conf/cvpr/HeCXLDG22) proposed Masked Autoencoders (MAE), demonstrating that reconstructing a high proportion of masked image patches using an asymmetric encoder-decoder architecture can achieve strong visual representations, rivaling or surpassing supervised pre-training on downstream tasks. Subsequently, Huang et al. (DBLP:conf/nips/HuangCGXZLLX22) introduced MaskedGAN, which uses randomized spatial and frequency masking to stabilize GAN training under data-scarce conditions, improving the robustness of generative models. Moving from representation learning to the domain of out-of-distribution detection, Li et al. (DBLP:journals/corr/abs-2302-02615) argued that reconstruction-based objectives are inherently better at learning in-distribution representations and thus more effective at detecting out-of-distribution samples, outperforming recognition-based methods. Xie et al. (DBLP:conf/iclr/0002LZ0OL23) extended MIM to the frequency domain with Masked Frequency Modeling (MFM), where the model learns to predict masked frequency components instead of spatial patches, capturing structured global image priors more efficiently. In the context of deepfake detection, Das et al. (Das2023UnmaskingDM) leveraged masked autoencoding with spatiotemporal transformers for videos, showing that masked modeling can facilitate strong generalization across datasets by encoding both spatial and temporal inconsistencies in fake content. More recently, Chen et al. (Chen2024MaskedCD) applied masking in conjunction with a conditional diffusion model to augment training data, demonstrating improved generalization to unseen deepfake forgeries. Building on these insights, our method incorporates random masking directly into supervised training, using frequency-domain masking to enforce feature robustness while maintaining low computational cost.
@@ 2.4. Efficient and Green AI
The increasing scale and complexity of deep neural networks have led to rising concerns over energy consumption and deployment cost, sparking interest in efficient and Green AI methods. Wen et al. (Wen2016LearningSS) addressed this early by introducing Structured Sparsity Learning (SSL), a method that enforces group-wise sparsity in network parameters to enable hardware-friendly acceleration without degrading performance. Liu et al. (Liu2017LearningEC) proposed Network Slimming, an approach that prunes unimportant channels via sparsity-induced scaling factors during training, achieving compact models compatible with standard hardware. Wang et al. (Wang2020NeuralPV) further advanced pruning techniques by introducing growing regularization to facilitate neural network compression without requiring expensive Hessian-based computations. Lee et al. (Lee2020LayeradaptiveSF) built on this with a layer-adaptive sparsity mechanism that assigns optimal sparsity levels across layers using a magnitude-based scoring strategy, improving the sparsity-performance tradeoff without manual tuning. Chen et al. (Chen2022DefakeHopAE) developed DefakeHop++, a lightweight deepfake detector that improves detection accuracy through expanded facial region coverage and supervised feature selection, outperforming CNN-based counterparts while maintaining a low parameter count. In a similar study, Vidya et al. (K2023CompressedDD) presented a spatio-temporal deepfake detector for compressed videos, employing model pruning and feature fusion to retain high performance with reduced model complexity. Fang et al. (Fang2023DepGraphTA) proposed DepGraph, a general structural pruning framework that models inter-layer dependencies to enable consistent and automatic pruning across various architectures, eliminating the need for manual design. Most recently, Lim et al. (Lim2024DistilDIREAS) proposed DistilDIRE, a diffusion-aware, lightweight deepfake detector that distills knowledge from large models to produce compact, fast models capable of accurate detection in real-time scenarios. Collectively, these works exemplify the trend towards efficient and scalable AI, aligning closely with our own method that emphasizes frequency-domain masking as a lightweight and pruning-resilient training strategy for sustainable universal deepfake detection (Kuo2022GreenLI; Zhu2021APixelHopAG; Zhu2022RGGIDAR; chen2024gift; Wei2024AGL).
@@ 4. Experiments
Datasets: Our experiments adhere to the training/validation setup of Wang et al. (wang2019cnngenerated), using ProGAN (720k training and 4k validation samples). For testing, we evaluate on the following datasets:
@@
  • Wang et al. (wang2019cnngenerated): 11 datasets covering GANs (ProGAN, StyleGAN, BigGAN, GauGAN, CycleGAN, StarGAN), perceptual-loss models (CRN, IMLE), low-level vision (SITD, SAN), and DeepFake images.
@@
  • Ojha et al. (ojha2023fakedetect): 8 diffusion model datasets including Guided Diffusion, Latent Diffusion (LDM) with varying steps of noise refinement (e.g., 100, 200) and generation guidance (w/ CFG), Glide which used two stages of noise refinement steps. The first stage (e.g., 50, 100) is to get low-resolution 64 × 64 images and then use 27 steps to upsample the image to 256 × 256 in the next stage, and lastly DALL-E-mini.
@@
  • Das et al. (Das2024BDfreshwaterfishAI): 2 synthetic fish datasets (ControlNet and Stable Diffusion) at 512x512 resolution with neutral backgrounds, avoiding underwater scenes.
@@
State-of-the-art (SOTA): We compared with following methods: Chai et al. (DBLP:conf/eccv/ChaiBLI20), Natraj et al. (Nataraj2019DetectingGG), Zhang et al. (Zhang2019DetectingAS), Ojha et al. (ojha2023fakedetect), Tan et al. (Tan2024RethinkingTU), Wang et al. (wang2019cnngenerated) and Gragnaniello et al. (Gragnaniello2021). Wang et al. (wang2019cnngenerated) and Gragnaniello et al. (Gragnaniello2021) methods achieved SOTA detection performance across GANs and diffusion models according to (Corvi_2023_ICASSP). We applied our proposed frequency masking on these SOTA and evaluated the improvement.
@@
Data processing: During training, we simulate various image post-processing operations by training all models with randomly left-right flipped and 224-pixel cropped images, while evaluating several augmentation variants: (1) Blur+JPEG (0.5): 50% probability for Gaussian blur and JPEG, and (2) Blur+JPEG (0.1): same as (1) but with 10% probability. These augmentation strategies help improve model robustness against common image manipulations encountered in real-world scenarios. The specific probabilities were chosen based on the work of Wang et al. (wang2019cnngenerated) showing their effectiveness in balancing regularization strength and preservation of original image features.
@@
Evaluation: Following other works (wang2019cnngenerated; ojha2023fakedetect; Gragnaniello2021), we evaluate our model’s performance on each dataset separately using mean average precision (mAP) and Area Under Receiver Operating Curve (AUROC), as these threshold-less, ranking-based scores are not sensitive to the threshold of real versus fake images and we expect performance to depend on each dataset’s semantic content. During testing, each image is center-cropped to $\times224$ pixels without resizing to match the training post-processing pipeline, with no data augmentation applied. This consistent evaluation protocol ensures fair comparison across different models and datasets. Additionally, we report both metrics to provide comprehensive insights into model performance across different operational points.
@@
As our default setting, we used Wang et al. (wang2019cnngenerated) and incorporated image augmentations like Gaussian Blur and JPEG compression, each having a 10% likelihood of being applied. This configuration was selected as it provided the best trade-off between generalization performance and computational capacity in our experiments.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: TOMMccs: Computing methodologies Biometrics
@@ 1. Introduction
The proliferation of increasingly convincing synthetic images, facilitated by generative AI, poses significant challenges across multiple sectors, including cybersecurity, digital forensics, and public discourse (Ojha et al., 2023; Gragnaniello et al., 2021; Chai et al., 2020; Abdollahzadeh et al., 2023). These AI-generated images could be misused as deepfakes for malicious purposes, such as disinformation. Because these fake images can look very real, they can trick people and systems easily. Detection of deepfakes is an important problem that has attracted significant attention.
@@
Universal deepfake detection. Early research in deepfake detection primarily focused on identifying synthetic images generated by specific types of generative models, such as Generative Adversarial Networks (GANs) (Mirsky and Lee, 2020). However, with the rapid advancement of generative methods, including the rise of diffusion models (Rombach et al., 2021a), the need for universal deepfake detection has become more prominent. This approach aims to detect deepfakes across a wide variety of generative AI models, including those that were not part of the training data for the detection system. Such a shift highlights the necessity for deepfake detection systems to exhibit robust generalization capabilities. In the early stages, Wang et al. (Wang et al., 2020b) explored the use of post-processing and data augmentation techniques to detect synthetic images, focusing particularly on those generated by GANs. Later, Chandrasegaran et al. (Chandrasegaran et al., 2022) introduced the concept of transferable forensic features, which are generalizable across different generative models, thus improving the detector’s ability to handle unseen models. More recent approaches, such as those by Ojha et al. (Ojha et al., 2023), leverage the feature space of large pretrained models, showing strong performance across unseen generators, although they depend heavily on access to large-scale pretrained models. A notable advancement by Tan et al. (Tan et al., 2024a) investigated the impact of architectural artifacts, specifically those introduced by upsampling operations in generative models, and proposed detecting structural inconsistencies to improve generalization. Lastly, Chen et al. (Chen et al., 2022b) examined one-shot test-time training to enhance model generalization; however, this technique introduces a trade-off between accuracy and computational efficiency due to the need for additional resources per test sample.
@@
The case for Green AI. Universal deepfake detection, which aims for strong generalization across a variety of scenarios, often depends on large pre-trained models (Ojha et al., 2023) and specialized feature extraction techniques (Tan et al., 2024b), both of which demand substantial computational resources. This dependency raises concerns about the scalability of these methods and their negative environmental impact due to the high energy consumption associated with training and inference. In response, the machine learning community is increasingly focusing on the concept of Green AI, which promotes the development of methods capable of achieving competitive performance while minimizing energy usage and computational cost (Chen et al., 2024a; Kuo and Madni, 2022; Zhu et al., 2022, 2021; Wei et al., 2024). This shift in focus has led to efforts aimed at creating lightweight yet generalizable detection strategies that remain effective even in environments with limited resources. In this regard, frequency-domain representations stand out as an especially promising solution (Coccomini et al., 2024). These representations preserve important structural and statistical information, typically with fewer dimensions than their spatial counterparts, and require significantly lower computational overhead. Consequently, employing a frequency-based approach offers the potential to enhance both the generalization ability and efficiency of deepfake detection models. Beyond general-purpose detection, specialized domains (e.g., aquaculture) increasingly rely on synthetic data for training and augmentation. However, poor-quality synthetic samples may propagate errors in critical applications—from misdiagnosing fish diseases to inaccurate medical assessments. A robust universal detector must therefore generalize not only across generative architectures but also across diverse application domains with potentially limited training data.
@@
Masked image modeling. In recent years, masked image modeling (MIM) has gained significant attention as an effective self-supervised pre-training technique to enhance the generalization ability of models (He et al., 2022; Li et al., 2023; Xie et al., 2023; Huang et al., 2022). This method revolves around pre-training models to predict missing or masked parts of unlabeled data, with reconstruction loss serving as the objective for learning. After the pre-training phase, the trained encoder can be effectively fine-tuned for a variety of downstream tasks, demonstrating strong transferability. He et al. (He et al., 2022) demonstrated that training with a masked autoencoder (MAE) allows high-capacity models to achieve state-of-the-art performance in terms of generalization across diverse applications. More recent works, such as those by Li et al. (Li et al., 2023), argue that reconstruction-based objectives excel at learning in-distribution representations, making them particularly effective for out-of-distribution detection. Xie et al. (Xie et al., 2023) expanded this approach by incorporating frequency-domain information into the masking process, proposing Masked Frequency Modeling (MFM), where the model learns to predict masked frequency components rather than spatial patches. This enables models to capture structured global image priors more efficiently. In the context of deepfake detection, Das et al. (Das et al., 2023) leveraged spatiotemporal transformers and masked autoencoding, showing that incorporating both spatial and temporal inconsistencies improves generalization across datasets.
@@
We analyzed both random masking and geometric transformations for universal deepfake detection. Our results suggest that frequency masking is more effective in generalizing detection across diverse generative AI approaches. Our finding is consistent with a recent study by Corvi et al. (Corvi et al., 2023a), which identifies frequency artifacts in GAN- and diffusion-based synthetic images. Different from (Corvi et al., 2023a), our main contribution is a new training method that improves detection accuracy and robustness via frequency-domain masking. We further demonstrate that this method is resilient to model compression, i.e., structured pruning, making it attractive for real-world, resource-constrained deployments.
We remark that most existing detectors focus on spatial-domain artifacts (Ojha et al., 2023; Gragnaniello et al., 2021; Wang et al., 2020b). Our contributions are summarized as follows:
@@ 2.1. Universal Deepfake Detection
Early research in universal deepfake detection focused on identifying common artifacts in synthetic images. Zhang et al. (Zhang et al., 2019) proposed simulating GAN artifacts to train frequency-domain classifiers capable of generalizing across unseen models. Wang et al. (Wang et al., 2020b) discovered that CNN-generated images exhibit systematic artifacts detectable by CNN-based classifiers, even when the detector is trained on a single generator. This finding suggested shared weaknesses among early GAN models. In contrast, Gragnaniello et al. (Gragnaniello et al., 2021) emphasized the growing realism of synthetic content and highlighted the necessity of robust and generalizable detection tools in the face of increasingly sophisticated generative models. Chandrasegaran et al. (Chandrasegaran et al., 2022) introduced the idea of transferable forensic features, demonstrating that certain features—particularly color-based ones—can be learned to generalize across models. Ojha et al. (Ojha et al., 2023) proposed leveraging the feature space of large pretrained models for real-vs-fake classification. This method, without explicit model-specific training, was shown to outperform traditional classifiers on unseen generative models. More recently, Tan et al. (Tan et al., 2024a) explored architectural artifacts, particularly those introduced by upsampling operations in generative models. They proposed detecting structural inconsistencies such as neighboring pixel relationships (NPR), which yielded high generalization performance on unseen generators. Compared to these approaches, our method introduces masked image modeling as a training-time augmentation that reduces overfitting to generator-specific traces.
@@ 2.2. Frequency-Domain Analysis
Pioneering studies have explored frequency-domain cues as critical features for improving the robustness and generalizability of deepfake detection systems. Frank et al. (Frank et al., 2020) first demonstrated that GAN-generated images often exhibit unnatural frequency artifacts due to upsampling operations, which can be effectively exploited to identify fakes. Qian et al. (Qian et al., 2020) further investigated these frequency-aware clues, showing that subtle forgery traces and compression-induced anomalies manifest more clearly in the frequency domain, especially in low-quality images. Luo et al. (Luo et al., 2021) addressed the generalization challenge across forgery methods by proposing three modules to extract high-frequency noise patterns, which generalize better than spatial-domain textures. Complementary to this, Le and Woo (Le and Woo, 2021) introduced a knowledge distillation framework where a student network learns from high-quality teacher signals in the frequency domain, enabling robust detection of compressed and degraded images. Jeong et al. (Jeong et al., 2022) recognized the risk of overfitting to frequency-specific artifacts and proposed FrePGAN, which transitions from frequency-focused learning to incorporating image-level irregularities, thus balancing generalization across known and unknown generators. Tian et al. (Tian et al., 2023) combined RGB and frequency representations using attention-based fusion, showing that this dual-modality approach alleviates overfitting and enhances generalization. Most recently, Tan et al. (Tan et al., 2024b) developed a frequency-aware architecture that targets high-frequency patterns across spatial and channel dimensions to improve generalizability while maintaining low model complexity. Li et al. (Li et al., 2024) introduced a data generation strategy that creates pseudo-fake images by blending frequency-domain features, enriching training data to better capture generalizable forgery traits. Distinctly, our approach employs frequency-domain masking as a supervised augmentation method that promotes learning of features resilient to domain shift.
@@ 2.3. Masked Image Modeling
Masked image modeling (MIM) has emerged as a powerful self-supervised learning paradigm in computer vision, with increasing relevance to generalization and robustness. He et al. (He et al., 2022) proposed Masked Autoencoders (MAE), demonstrating that reconstructing a high proportion of masked image patches using an asymmetric encoder-decoder architecture can achieve strong visual representations, rivaling or surpassing supervised pre-training on downstream tasks. Subsequently, Huang et al. (Huang et al., 2022) introduced MaskedGAN, which uses randomized spatial and frequency masking to stabilize GAN training under data-scarce conditions, improving the robustness of generative models. Moving from representation learning to the domain of out-of-distribution detection, Li et al. (Li et al., 2023) argued that reconstruction-based objectives are inherently better at learning in-distribution representations and thus more effective at detecting out-of-distribution samples, outperforming recognition-based methods. Xie et al. (Xie et al., 2023) extended MIM to the frequency domain with Masked Frequency Modeling (MFM), where the model learns to predict masked frequency components instead of spatial patches, capturing structured global image priors more efficiently. In the context of deepfake detection, Das et al. (Das et al., 2023) leveraged masked autoencoding with spatiotemporal transformers for videos, showing that masked modeling can facilitate strong generalization across datasets by encoding both spatial and temporal inconsistencies in fake content. More recently, Chen et al. (Chen et al., 2024b) applied masking in conjunction with a conditional diffusion model to augment training data, demonstrating improved generalization to unseen deepfake forgeries. Building on these insights, our method incorporates random masking directly into supervised training, using frequency-domain masking to enforce feature robustness while maintaining low computational cost.
@@ 2.4. Efficient and Green AI
The increasing scale and complexity of deep neural networks have led to rising concerns over energy consumption and deployment cost, sparking interest in efficient and Green AI methods. Wen et al. (Wen et al., 2016) addressed this early by introducing Structured Sparsity Learning (SSL), a method that enforces group-wise sparsity in network parameters to enable hardware-friendly acceleration without degrading performance. Liu et al. (Liu et al., 2017) proposed Network Slimming, an approach that prunes unimportant channels via sparsity-induced scaling factors during training, achieving compact models compatible with standard hardware. Wang et al. (Wang et al., 2020a) further advanced pruning techniques by introducing growing regularization to facilitate neural network compression without requiring expensive Hessian-based computations. Lee et al. (Lee et al., 2020) built on this with a layer-adaptive sparsity mechanism that assigns optimal sparsity levels across layers using a magnitude-based scoring strategy, improving the sparsity-performance tradeoff without manual tuning. Chen et al. (Chen et al., 2022a) developed DefakeHop++, a lightweight deepfake detector that improves detection accuracy through expanded facial region coverage and supervised feature selection, outperforming CNN-based counterparts while maintaining a low parameter count. In a similar study, Vidya et al. (K et al., 2023) presented a spatio-temporal deepfake detector for compressed videos, employing model pruning and feature fusion to retain high performance with reduced model complexity. Fang et al. (Fang et al., 2023) proposed DepGraph, a general structural pruning framework that models inter-layer dependencies to enable consistent and automatic pruning across various architectures, eliminating the need for manual design. Most recently, Lim et al. (Lim et al., 2024) proposed DistilDIRE, a diffusion-aware, lightweight deepfake detector that distills knowledge from large models to produce compact, fast models capable of accurate detection in real-time scenarios. Collectively, these works exemplify the trend towards efficient and scalable AI, aligning closely with our own method that emphasizes frequency-domain masking as a lightweight and pruning-resilient training strategy for sustainable universal deepfake detection (Kuo and Madni, 2022; Zhu et al., 2021, 2022; Chen et al., 2024a; Wei et al., 2024).
@@ 4. Experiments
Datasets: Our experiments adhere to the training/validation setup of Wang et al. (Wang et al., 2020b), using ProGAN (720k training and 4k validation samples). For testing, we evaluate on the following datasets:
@@
  • Wang et al. (Wang et al., 2020b): 11 datasets covering GANs (ProGAN, StyleGAN, BigGAN, GauGAN, CycleGAN, StarGAN), perceptual-loss models (CRN, IMLE), low-level vision (SITD, SAN), and DeepFake images.
@@
  • Ojha et al. (Ojha et al., 2023): 8 diffusion model datasets including Guided Diffusion, Latent Diffusion (LDM) with varying steps of noise refinement (e.g., 100, 200) and generation guidance (w/ CFG), Glide which used two stages of noise refinement steps. The first stage (e.g., 50, 100) is to get low-resolution 64 × 64 images and then use 27 steps to upsample the image to 256 × 256 in the next stage, and lastly DALL-E-mini.
@@
  • Das et al. (Das et al., 2024): 2 synthetic fish datasets (ControlNet and Stable Diffusion) at 512x512 resolution with neutral backgrounds, avoiding underwater scenes.
@@
State-of-the-art (SOTA): We compared with following methods: Chai et al. (Chai et al., 2020), Natraj et al. (Nataraj et al., 2019), Zhang et al. (Zhang et al., 2019), Ojha et al. (Ojha et al., 2023), Tan et al. (Tan et al., 2024a), Wang et al. (Wang et al., 2020b) and Gragnaniello et al. (Gragnaniello et al., 2021). Wang et al. (Wang et al., 2020b) and Gragnaniello et al. (Gragnaniello et al., 2021) methods achieved SOTA detection performance across GANs and diffusion models according to (Corvi et al., 2023b). We applied our proposed frequency masking on these SOTA and evaluated the improvement.
@@
Data processing: During training, we simulate various image post-processing operations by training all models with randomly left-right flipped and 224-pixel cropped images, while evaluating several augmentation variants: (1) Blur+JPEG (0.5): 50% probability for Gaussian blur and JPEG, and (2) Blur+JPEG (0.1): same as (1) but with 10% probability. These augmentation strategies help improve model robustness against common image manipulations encountered in real-world scenarios. The specific probabilities were chosen based on the work of Wang et al. (Wang et al., 2020b) showing their effectiveness in balancing regularization strength and preservation of original image features.
@@
Evaluation: Following other works (Wang et al., 2020b; Ojha et al., 2023; Gragnaniello et al., 2021), we evaluate our model’s performance on each dataset separately using mean average precision (mAP) and Area Under Receiver Operating Curve (AUROC), as these threshold-less, ranking-based scores are not sensitive to the threshold of real versus fake images and we expect performance to depend on each dataset’s semantic content. During testing, each image is center-cropped to $\times224$ pixels without resizing to match the training post-processing pipeline, with no data augmentation applied. This consistent evaluation protocol ensures fair comparison across different models and datasets. Additionally, we report both metrics to provide comprehensive insights into model performance across different operational points.
@@
As our default setting, we used Wang et al. (Wang et al., 2020b) and incorporated image augmentations like Gaussian Blur and JPEG compression, each having a 10% likelihood of being applied. This configuration was selected as it provided the best trade-off between generalization performance and computational capacity in our experiments.